The following table shows the amount of converted sugar in a chemical process at different temperatures.
Temperature, x Converted Sugar, y
1.2 8.2
1.4 8.5
1.6 8.4
1.8 9.3
2.0 8.9
2.2 10.5
2.4 9.3


a. Compute for the correlation coefficient
b. Estimate the linear regression line.
c. Estimate the mean amount of sugar produced when the temperature recorded is 1.7.
1
Expert's answer
2020-10-19T17:05:29-0400

Solution

a. Correlation coefficient, r.


"r={n \sum (xy)- (\sum x)( \sum y) \over \sqrt{[n\sum x^2 - ({\sum x})^2][n\sum y^2 - ({\sum y})^2]}}""={(7*115.14)-(12.6*63.1) \over \sqrt{[7*23.8 - {(12.6)}^2][7*572.49-{(63.1)}^2]}}""={10.92 \over 14.2277} = 0.76751"

Ans: 0.76751


b. Linear regression line

"y=a+bx"


"b= {n \sum (xy)- (\sum x)( \sum y) \over [n\sum x^2 - ({\sum x})^2]}""={7*115.14 - (12.6*63.1) \over (7*23.8 - {(12.6)}^2} = {10.92 \over 7.84 }""=1.3929"

"a= {\sum y - b\sum x \over n}""={63.1 - 1.3929(12.6) \over 7}""=6.5071"

"\therefore y=6.5071+1.3929x"


c. mean amount of sugar when temp. is 1.7


"y=6.5071+1.3929(1.7)""=8.87503"